# Equivariant scaling asymptotics for Poisson and Szegő kernels on Grauert tube boundaries

### v1

@@ 6 Near-graph uniform asymptotic expansions
The first statement is obvious by definition. If the second was false, for any $j=1,2,\ldots$ there would exist $(y_{j}^{\prime},y_{j}^{\prime\prime})\in Z^{\tau}\times Z^{\tau}$ having distance $<1/j$ from $(x_{1},x_{2})$ and $(g_{j},t_{j})\in\Sigma_{\chi}(y_{j}^{\prime},y_{j}^{\prime\prime})$ having distance from $\Sigma_{\chi}(x_{1},x_{2})$ no less than $\epsilon_{0}$, for some fixed $\epsilon_{0}>0$. By compactness, we may assume without loss that $g_{j}\rightarrow g_{\infty}\in G$ and $t_{j}\rightarrow t_{\infty}$ in $\mathrm{supp}(\chi)$. By continuity,
@@
::=:= $\Xi_{\nu}\left(h_{1}^{-1}\,\kappa_{l}^{-1}\,e^{-\xi}\right)\,\chi\left(t+t_{1}\right)\cdot\mathcal{V}\left(\theta,\mathbf{u}\right)\cdot\gamma(t,\theta,\xi,\mathbf{u})\cdot f_{1}(v)\cdot f_{2}(u)$
@@
$S$ ::=:= ${D_{x}^{l}}^{\dagger}\,{D_{x_{1}}^{l}}-{D_{x_{1}}^{l}}^{\dagger}\,(I_{2d-2}-\imath\,J_{0})\,\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}\,(I_{2d-2}+\imath\,J_{0})\,{D_{x_{1}}^{l}}.$
@@
$\Re(S)$ ::=:= ${D_{x_{1}}^{l}}^{\dagger}\,{D_{x}^{l}}-{D_{x}^{l}}^{\dagger}\,\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}\,\,{D_{x}^{l}}$
@@
$\Re(S)$ ::=:= ${D_{x_{1}}^{l}}^{\dagger}\,{D_{x_{1}}^{l}}-{D_{x_{1}}^{l}}^{\dagger}\,\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}\,\,{D_{x_{1}}^{l}}$
@@
where $\tilde{P}_{k+3j+a}$ is homogeneous of total degree $3j+a+k$. We obtain the same conclusions as before, that is general terms of the form $\lambda^{-k/2}\,P_{r}$ with $r\leq 3\,k$ and $r-k$ even.
@@ 7.1 The equivariant Weyl law for $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$
is a diffeomorphism onto a neighbourhood of $x$ in $X^{\tau}$ (here we use normal Heisenberg coordinates at each $x^{\prime}$ to identify $\mathbb{R}^{d_{G}}\cong\mathbb{R}^{d_{G}}_{t}\cong T_{x^{\prime}}^{t}X^{\tau}$). If we pass to the rescaled local parametrization
$x_{\lambda}(\mathbf{v}^{t}):=x+\frac{\mathbf{v}^{t}}{\sqrt{\lambda}}\qquad(x^{\prime}\in Z^{\prime},\,\mathbf{v}^{t}\in\mathbb{R}^{d_{G}}),$ (164)
@@
where $F_{k,l,\nu}(x;\cdot)$ is a polynomial of degree $\leq 3\,k$ and parity $k$.
@@
$\sim$ $\int_{Z^{\tau}}\,\mathrm{d}V_{Z^{\tau}}(x)\left[\int_{\mathbb{R}^{d_{G}}}\,\Pi^{\tau}_{\chi,\nu,\lambda}(x+\mathbf{v}^{t},x+\mathbf{v}^{t})\,\,\mathcal{V}(x+\mathbf{v}^{t})\,\mathrm{d}\mathbf{v}^{t}\right]$
$=$ $\lambda^{-d_{G}/2}\,\int_{Z^{\tau}}\,\mathrm{d}V_{Z^{\tau}}(x)\left[\int_{\mathbb{R}^{d_{G}}}\,\Pi^{\tau}_{\chi,\nu,\lambda}\left(x_{\lambda}(\mathbf{v}^{t}),x_{\lambda}(\mathbf{v}^{t})\right)\,\,\mathcal{V}(x_{\lambda}(\mathbf{v}^{t}))\,\mathrm{d}\mathbf{v}^{t}\right]$
where $\mathrm{d}V_{Z^{\tau}}$ is the Riemannian volume form on $Z^{\tau}$ for the restricted metric, and $\mathcal{V}(x_{\lambda}(\mathbf{0}))=1$ for every $x\in X^{\tau}$. Integration in $\mathrm{d}\mathbf{v}^{t}$ is now over an expanding ball of radius $O\left(\lambda^{\epsilon^{\prime}}\right)$ in $\mathbb{R}^{d_{G}}$. By multiplying the asymptotic expansion of Theorem 28 with the Taylor expansion of $\mathcal{V}(x_{\lambda}(\mathbf{v}^{t}))$, we obtain an asymptotic expansion for the integrand in (7.1) of the form
@@
$\sim$ $\frac{1}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\,\lambda^{-d_{G}/2}\cdot\dim(\nu)^{2}$
@@
$=$ $\frac{1}{2^{d-1}\,\sqrt{2\,\pi}}\,\left(\frac{\lambda}{\pi\,\tau}\right)^{d-1-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)\cdot\left[\chi(0)+\sum_{k\geq 1}\,\lambda^{-k}\,\mathfrak{f}_{k,\nu}\right].$
@@
$=$ $\frac{1}{2^{d-1}\,\sqrt{2\,\pi}}\,\cdot\frac{\lambda}{d-d_{G}}\,\left(\frac{\lambda}{\pi\,\tau}\right)^{d-1-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@
$\mathcal{W}_{\nu}^{\tau}(\lambda)$ $=$ $\frac{1}{2^{d-1}\,(2\,\pi)}\,\cdot\frac{\lambda}{d-d_{G}}\,\left(\frac{\lambda}{\pi\,\tau}\right)^{d-1-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@
$=$ $\frac{1}{2^{d_{G}}}\cdot\frac{\tau}{d-d_{G}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@ 7.4 Proof of Proposition 38 and Theorem 40
$\sim$ $\frac{1}{\sqrt{2\,\pi}}\,\left(\frac{1}{2\,\pi}\right)^{d-1-d_{G}/2}\left(\frac{\lambda}{\tau}\right)^{(d-1-d_{G})/2}\cdot\frac{\dim(\nu)^{2}}{V_{eff}(x_{1})}$
@@
$=$ $\frac{\pi^{d_{G}}\,\sqrt{2\,\pi}}{(2\,\pi)^{d}}\,\left(\frac{\lambda}{\tau}\right)^{\frac{d-1}{2}-d_{G}}\,\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@ 8 Index of Notation
    the matrices $A_{c}$, $P$, $Q$ associated to a symplectic matrix $A$: see Definition 8; the function $\Psi_{A}:\mathbb{R}^{2d-2}\times\mathbb{R}^{2d-2}\rightarrow\mathbb{C}$ associated to a symplectic matrix $A$: see Definition 9;
@@
    $A_{\chi}$: see Definition 22; $\mathcal{F}_{\chi,\lambda}$, $\mathcal{B}_{\nu}$: see Definition 23;

### v3

@@ 6 Near-graph uniform asymptotic expansions
The first statement is obvious by definition. If the second was false, for any $j=1,2,\ldots$ there would exist $(y_{j}^{\prime},y_{j}^{\prime\prime})\in Z^{\tau}\times Z^{\tau}$ having distance <1/j<1/j from $(x_{1},x_{2})$ and $(g_{j},t_{j})\in\Sigma_{\chi}(y_{j}^{\prime},y_{j}^{\prime\prime})$ having distance from $\Sigma_{\chi}(x_{1},x_{2})$ no less than $\epsilon_{0}$, for some fixed $\epsilon_{0}>0$. By compactness, we may assume without loss that $g_{j}\rightarrow g_{\infty}\in G$ and $t_{j}\rightarrow t_{\infty}$ in $\mathrm{supp}(\chi)$. By continuity,
@@
$:=$ $\Xi_{\nu}\left(h_{1}^{-1}\,\kappa_{l}^{-1}\,e^{-\xi}\right)\,\chi\left(t+t_{1}\right)\cdot\mathcal{V}\left(\theta,\mathbf{u}\right)\cdot\gamma(t,\theta,\xi,\mathbf{u})\cdot f_{1}(v)\cdot f_{2}(u)$
@@
$S$ $:=$ ${D_{x}^{l}}^{\dagger}\,{D_{x_{1}}^{l}}-{D_{x_{1}}^{l}}^{\dagger}\,(I_{2d-2}-\imath\,J_{0})\,\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}\,(I_{2d-2}+\imath\,J_{0})\,{D_{x_{1}}^{l}}.$
@@
$\Re(S)$ $:=$ ${D_{x_{1}}^{l}}^{\dagger}\,{D_{x}^{l}}-{D_{x}^{l}}^{\dagger}\,\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}\,\,{D_{x}^{l}}$
@@
$\Re(S)$ $:=$ ${D_{x_{1}}^{l}}^{\dagger}\,{D_{x_{1}}^{l}}-{D_{x_{1}}^{l}}^{\dagger}\,\left(I_{2d-2}+B^{\dagger}\,B\right)^{-1}\,\,{D_{x_{1}}^{l}}$
@@
where $\tilde{P}_{k+3j+a}$ is homogeneous of total degree $3j+a+k$. We obtain the same conclusions as before, that is general terms of the form $\lambda^{-k/2}\,P_{r}$ with $r\leq 3\,k$ and rkr-k even.
@@ 7.1 The equivariant Weyl law for $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$
is a diffeomorphism onto a neighbourhood of $x$ in $X^{\tau}$ (here we use normal Heisenberg coordinates at each $x^{\prime}$ to identify $\mathbb{R}^{d_{G}}\cong\mathbb{R}^{d_{G}}_{t}\cong T_{x^{\prime}}^{t}X^{\tau}$). In view of Corollary 35 of [P-2024], if $x=x^{\prime}+\mathbf{v}^{t}$ then
$\mathrm{d}V_{X^{\tau}}(x)=2^{d_{G}/2}\,\mathcal{V}(x^{\prime}+\mathbf{v}^{t})\,\mathrm{d}V_{Z^{\tau}}(x^{\prime})\,\mathrm{d}\mathbf{v}^{t},$ (177)
where $\mathrm{d}V_{Z^{\tau}}$ is the Riemannian volume form on $Z^{\tau}$ for the restricted metric, and $\mathcal{V}(x^{\prime})=1$ for any $x^{\prime}\in Z^{\prime}$.
If we pass to the rescaled local parametrization
$x_{\lambda}(\mathbf{v}^{t}):=x^{\prime}+\frac{\mathbf{v}^{t}}{\sqrt{\lambda}}\qquad(x^{\prime}\in Z^{\prime},\,\mathbf{v}^{t}\in\mathbb{R}^{d_{G}}),$ (178)
@@
$\cdot e^{-\frac{2}{\tau}\,\left\|\mathbf{v}^{t}\right\|^{2}}\cdot\left[\chi(0)+\sum_{k\geq 1}\,\lambda^{-k/2}\,F_{k,\chi,\nu}\left(x^{\prime};\mathbf{v}^{t}\right)\right],$
where $F_{k,\chi,\nu}(x^{\prime};\cdot)$ is a polynomial of degree $\leq 3\,k$ and parity $k$.
@@
$\sim$ $2^{d_{G}/2}\int_{Z^{\tau}}\,\mathrm{d}V_{Z^{\tau}}(x)\left[\int_{\mathbb{R}^{d_{G}}}\,\Pi^{\tau}_{\chi,\nu,\lambda}(x+\mathbf{v}^{t},x+\mathbf{v}^{t})\,\,\mathcal{V}(x+\mathbf{v}^{t})\,\mathrm{d}\mathbf{v}^{t}\right]$
$=$ $\left(\frac{\lambda}{2}\right)^{-d_{G}/2}\,\int_{Z^{\tau}}\,\mathrm{d}V_{Z^{\tau}}(x)\left[\int_{\mathbb{R}^{d_{G}}}\,\Pi^{\tau}_{\chi,\nu,\lambda}\left(x_{\lambda}(\mathbf{v}^{t}),x_{\lambda}(\mathbf{v}^{t})\right)\,\,\mathcal{V}(x_{\lambda}(\mathbf{v}^{t}))\,\mathrm{d}\mathbf{v}^{t}\right].$
Integration in $\mathrm{d}\mathbf{v}^{t}$ is now over an expanding ball of radius $O\left(\lambda^{\epsilon^{\prime}}\right)$ in $\mathbb{R}^{d_{G}}$. By multiplying the asymptotic expansion of Theorem 28 with the Taylor expansion of $\mathcal{V}(x_{\lambda}(\mathbf{v}^{t}))$, we obtain an asymptotic expansion for the integrand in (7.1) of the form
@@
$\sim$ $\frac{2^{d_{G}/2}}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\,\lambda^{-d_{G}/2}\cdot\dim(\nu)^{2}$
@@
$=$ $\frac{1}{2^{d_{G}/2}\,\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)\cdot\left[\chi(0)+\sum_{k\geq 1}\,\lambda^{-k}\,\mathfrak{f}_{k,\nu}\right].$
@@
$=$ $\frac{1}{2^{d_{G}/2}\,\sqrt{2\,\pi}}\,\cdot\frac{\lambda}{d-d_{G}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@
$\mathcal{W}_{\nu}^{\tau}(\lambda)$ $=$ $\frac{2^{d_{G}/2}}{2^{d-1}\,(2\,\pi)}\,\cdot\frac{\lambda}{d-d_{G}}\,\left(\frac{\lambda}{\pi\,\tau}\right)^{d-1-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@
$=$ $\frac{1}{2^{d_{G}/2}}\cdot\frac{\tau}{d-d_{G}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@ 7.4 Proof of Proposition 38 and Theorem 40
$\sim$ $\frac{1}{\sqrt{2\,\pi}}\,\left(\frac{1}{2}\right)^{d-1-d_{G}/2}\left(\frac{\lambda}{\pi\,\tau}\right)^{(d-1-d_{G})/2}\cdot\frac{\dim(\nu)^{2}}{V_{eff}(x_{1})}$
@@
$=$ $\frac{1}{\sqrt{2\pi}}\,\frac{1}{2^{d-1-d_{G}/2}}\,\left(\frac{\lambda}{\pi\,\tau}\right)^{\frac{d-1}{2}-d_{G}}\,\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@ 8 Index of Notation
    the matrices $A_{c}$, $P$, $Q$ associated to a symplectic matrix AA: see Definition 8; the function $\Psi_{A}:\mathbb{R}^{2d-2}\times\mathbb{R}^{2d-2}\rightarrow\mathbb{C}$ associated to a symplectic matrix AA: see Definition 9;
@@
    $A_{\chi}$: see Definition 22; $\mathcal{F}_{\chi}$, $\mathcal{B}_{\nu}$: see Definition 23;